# Can Large Language Models Be Trusted as Black-Box Evolutionary Optimizers for Combinatorial Problems?

### v1

@@ Abstract
Evolutionary computation excels in complex optimization but demands deep domain knowledge, restricting its accessibility. Large Language Models (LLMs) offer a game-changing solution with their extensive knowledge and could democratize the optimization paradigm. Although LLMs possess significant capabilities, they may not be universally effective, particularly since evolutionary optimization encompasses multiple stages. It is therefore imperative to evaluate the suitability of LLMs as evolutionary optimizer (EVO). Thus, we establish a series of rigid standards to thoroughly examine the fidelity of LLM-based EVO output in different stages of evolutionary optimization and then introduce a robust error-correction mechanism to mitigate the output uncertainty. Furthermore, we explore a cost-efficient method that directly operates on entire populations with excellent effectiveness in contrast to individual-level optimization. Through extensive experiments, we rigorously validate the performance of LLMs as operators targeted for combinatorial problems. Our findings provide critical insights and valuable observations, advancing the understanding and application of LLM-based optimization.
@@
Keywords Evolutionary optimization  $\cdot$ Large language models  $\cdot$ Combinatorial problems
1 Introduction
@@
The advancement of computing technology, especially large language models (LLMs) [1, 2], has made it possible to revolutionize traditional programming methodologies [3, 4, 5] and reduce human labor. This vision is already materializing in the field of optimization [6], where optimization of combinatorial problems is achieved not through mathematical analysis or traditional programming but via prompt engineering only. Evolutionary optimization has proven to be an exceptionally powerful technique for tackling complex problems due to its simplicity and effectiveness [7, 8]. Despite their benefits, implementing evolutionary optimization is labor-intensive and demands deep domain expertise for precise execution [9]. This challenge intensifies with constrained problems or when tailoring methods for specific needs. Such difficulty is especially pronounced for non-technical users, presenting a major barrier to accessibility and effective application. Therefore, we ask the critical question of whether LLMs can directly serve as evolutionary optimizer (EVO).
To date, some preliminary studies have explored the use of LLMs as crossover operators [10], mutation operators [11], and in combination [12]. Despite these advancements [13, 14], LLMs have shown variable performance across different domains; they excel in some areas [15, 16, 17] but not in tasks requiring precise arithmetic and logical reasoning [18, 19]. For most researchers, it is challenging to train a large model from scratch targeted to a specific problem or even fine-tune it due to the huge computational demands. Thus, when deploying these models for specific tasks, it is imperative to conduct a thorough analysis to ascertain whether LLMs are genuinely suitable for the task at hand.
@@
While these efforts have opened up LLMs as the black-box EVO, there is still a lack of comprehensive investigation to evaluate the effectiveness and reliability of LLMs as EVO, and to identify the factors affecting their performance. Huang et al. [20] conducted an investigation on LLMs as normal optimizers on different problems, but it is only targeted at one-shot optimization, and some important aspects such as reliability, scalability, and computational cost remain underexplored. Thus, our work here aims to serve as a thorough assessment of the performance of LLMs as operators in all stages of evolutionary optimization, some of which, like LLM-based initialization and selection, still remain underexplored in the existing literature. We aim to provide insights into the suitability of LLMs, clarifying their capabilities and limitations. Our evaluations compare LLM-based approaches with software-based evolutionary computation to assess the accuracy of LLMs in meeting expected outcomes and their effectiveness in fulfilling customized requirements.
@@
In addition, we explore a cost-effective method that utilises the entire population as optimization unit for the LLM-based reproduction operator and compare this method with another common approach that implements optimization on individual-by-individual. We specifically analyze the quality of generated solutions by these two distinct ways as well as their computational overhead. Some of our findings in this work are summarized as follows:
$\bullet$ LLMs can perform the selection, crossover, and mutation operations well. Their performance is relatively independent of dataset size and type, being more influenced by hyperparameters such as population size and solution size instead.
$\bullet$ LLMs may not be suitable for tasks involving intensive computation, such as initialization, and their performance tends to degrade greatly as the dataset size increases.
$\bullet$ The performance of LLMs relies on the amount of input data, and LLMs may encounter various issues during the evolutionary optimization process, making correction mechanisms indispensable for successful optimization.
2 LLM-based evolutionary optimizer
@@
As in the combinatorial problem within the evolutionary optimization framework, the solution representation is generally defined as
@@
where X depends on the specific task and $n$ is the solution size [24, 25]. In this work, we choose a node selection problem [26, 27] as the illustrative example. The problem is defined as: Given a graph $G=(V,E)$, where $V$ represents the set of nodes and $E$ represents the set of edges, the objective is to find a subset of nodes $S\subseteq V$ that maximizes the influence across the network. As for network-related problems, the solution representation can simply be
@@
Note that our focus is not on solving any specific network-related problem but on evaluating the performance of LLMs as operators for combinatorial problems. The reason for selecting this problem is that it is general enough to provide insights and observations for LLMs as EVO due to the representation similarity on combinatorial problems. To enhance generality, ‘element’ will replace ‘node’, and ‘dataset’ will replace ‘graph’ in the rest discussion. This generality is also reflected in our prompt engineering where little network domain knowledge is needed (please refer to the Appendix). For the selection, the inputs to LLMs is
@@
The first phase, Initialization is defined as $\mathcal{P}_{0}=\mathbf{F}_{I}(G)$, where $\mathcal{P}_{0}$ refers to the initialized population and $G$ refers to the input data. A strategic reduction in search space can effectively speed up convergence and improve effectiveness [28]. However, based on current research [18], it is not feasible to feed datasets directly to LLMs, either in the form of adjacency matrices or text descriptions, because the reasoning ability of current LLMs on graphs remains limited. Therefore, instead of providing LLMs with just a dataset, we opt to input element IDs along with easily accessible metrics. Following this, LLMs will be instructed to (1) rank the element based on a specific metric; (2) select a percentage of top elements as candidates; and (3) sample filtered elements to form the initialized population. Several basic restrictions must be followed to ensure the quality of the initialized population: (1) Each solution in the population should contain the same number of elements with no duplicates; (2) The solutions within the population should be as diverse as possible; and (3) elements outside the candidate set are not allowed to appear in the population.
@@
2.1 Computational cost of population- and individual-level LLM-based EVO
@@
[…]
@@
Notably, there are several restrictions that must be followed during reproduction: (1) The new solution should not contain duplicate/invalid elements; (2) The new solution should maintain the same size as the previous solution; and (3) The population size should remain the same as the previous population (only applicable to the population-level EVO).
@@
Table 1: The errors of LLM outputs during evolutionary optimization.
2.2 Error repair
@@
Table 2: The required validations for the LLM output in different phases of evolutionary optimization.
The probabilistic nature of LLMs can lead to occasional undesirable results. As evolutionary optimization is an iterative process, an error early in the process can ‘blow up’ in subsequent cycles, creating a cascading effect to ruin the optimization process. We therefore investigate the possible errors and set rigid standards to validate the LLM outputs in different phases. These errors are categorized as format errors and quality errors, with the latter further divided into critical errors and moderate errors. Specifically,

### v2

# Can Large Language Models Be Trusted as Evolutionary Optimizers for Network-Structured Combinatorial Problems?

@@ Abstract
Large Language Models (LLMs) have shown impressive capabilities in language understanding and reasoning across diverse domains. Recently, there has been increasing interests in utilizing LLMs not merely as assistants in optimization tasks, but as active optimizers, particularly for network-structured combinatorial problems. However, before LLMs can be reliably deployed in this role, a fundamental question must be addressed: Can LLMs iteratively manipulate solutions that consistently adhere to problem constraints? In this work, we propose a systematic framework to evaluate the capacity of LLMs to engage with problem structures. Rather than treating the model as a black-box generator, we adopt the commonly used evolutionary operators as optimizer and propose a comprehensive evaluation framework that rigorously assesses the output fidelity of LLM-generated operators across different stages of the evolutionary process. To enhance robustness, we introduce a hybrid error-correction mechanism that mitigates uncertainty in LLM outputs. Moreover, we develop a cost-efficient population-level optimization strategy that significantly improves efficiency compared to traditional individual-level approaches. Extensive experiments on a representative node-level combinatorial network optimization task demonstrate the effectiveness, adaptability, and inherent limitations of LLM-based operators. Our findings offer new perspectives on the integration of LLMs into evolutionary computation, providing practical insights for scalable optimization in networked systems.
Index Terms:
@@
Complex networks, combinatorial problems, large language models, evolutionary optimization.
I Introduction
@@
Large Language Models (LLMs), such as GPT (Generative Pre-trained Transformer) trained on vast datasets [1, 2, 3] have the capability to understand and generate human-like text based on the learned patterns. LLMs excel in tasks ranging from simple text completion to complex question answering, demonstrating a nuanced understanding of language context and semantics [4, 5]. The advancement of LLMs [6, 7], has made it possible to revolutionize traditional programming methodologies [8, 9] and reduce human labor. Traditional optimization is labor-intensive and demands domain expertise for precise execution [10]. This challenge intensifies with constrained problems or when tailoring methods for specific needs. Such difficulty is especially pronounced for non-technical users, presenting a major barrier to accessibility and effective application.
There is a growing interest in leveraging LLM not merely as auxiliary tools within optimization pipelines, but as primary solvers [11, 12]. This vision has already been materialized in [13], where optimization of combinatorial problems is achieved not through mathematical analysis or traditional programming but via prompt engineering only. In this work, some combinatorial problems have been used as illustrative examples, such as the traveling salesman problems (TSP) and the solution-score pair is fed into LLMs for the refined solution. Evolutionary optimization has long been recognized as a powerful approach for solving complex problems, valued for both its simplicity and effectiveness. Motivated by these strengths, recent efforts have begun to explore the use of LLMs within evolutionary optimization frameworks. For example, Liu et al. [14] proposed a framework called LMEA, making LLMs function as crossover and mutation operators to solve TSP with up to 20 nodes. Brahmachary et al. [15] leveraged LLMs to optimize the continuous problem with elitism mechanism and demonstrate the potential of LLM-based optimizer. Furthermore, liu et al. [16] utilized LLMs to serve as black-box optimizers for decomposition-based MOEA in a zero-shot manner. In [17], Meyerson et al. explored different structures that LLMs can operate on as crossover, including binary strings, mathematical expressions, text and code.
Despite these advancements, LLMs have shown variable performance across different domains; they excel in some areas such as strategy formulation [18, 19, 20] but not in tasks requiring precise arithmetic and logical reasoning [21, 22]. Therefore, for LLMs to reliably and robustly fulfill this role, a foundational challenge must be addressed: Can LLMs iteratively refine solutions while ensuring that all domain-specific constraints are preserved throughout the optimization process? In combinatorial settings, particularly those involving graphs, schedules, or other structured domains, maintaining feasibility is a challenging and essential requirement. Although encoding schemes differ considerably depending on the task, such as using permutation-based encodings for scheduling problems and binary or graph-based representations for network-related tasks, the fundamental process of manipulating these encodings remains consistent within the evolutionary algorithm framework. Core operations like crossover and mutation are applied universally, typically involving the exchange or replacement of elements, regardless of the specific encoding. Therefore, despite variations in representational strategies, evolutionary algorithms adopt a unified methodology for manipulate the solution space.
@@
To date, there is still a lack of comprehensive investigation to evaluate the effectiveness and reliability of LLMs as optimizers, and to identify the factors affecting their performance. Huang et al. [23] conducted an investigation on LLMs as normal optimizers on different problems, but it is only targeted at one-shot optimization, and some important aspects such as reliability, scalability, and computational cost remain underexplored. Due to the versatility of evolutionary optimizer on the network-related problem, it will serve the illustrative example in our work. Thus, we here aim to conduct a thorough assessment of the performance of LLMs as operators in all stages of evolutionary optimization, some of which, like LLM-based initialization and selection, still remain underexplored in the existing literature. We will also provide insights into the suitability of LLMs, clarifying their capabilities and limitations. Our evaluations compare LLM-based approaches with software-based evolutionary computation to assess the accuracy of LLMs in meeting expected outcomes and their effectiveness in fulfilling customized requirements.
@@
In addition, we explore a cost-effective method that utilizes the entire population as optimization unit for the LLM-based reproduction operator and compare this method with another common approach that implements optimization on individual-by-individual. We specifically analyze the quality of generated solutions by these two distinct ways as well as their computational overhead. Some of our findings in this work are summarized as follows:
$\bullet$ LLMs are capable of performing selection, crossover, and mutation operations within evolutionary optimization. However, their effectiveness is highly sensitive to hyperparameters such as population size and solution length, which must be carefully tuned to maintain performance.
$\bullet$ LLMs can effectively perform decision-making tasks such as selection, often outperforming traditional heuristic methods in adaptability. Unlike fixed-rule heuristics, LLMs can incorporate complex contextual information such as fitness values, element rankings when making selection decisions.
$\bullet$ The effectiveness and reliability of LLM-based evolutionary optimization are closely tied to the capacity of the underlying language model, particularly its number of parameters. Larger models generally demonstrate stronger reasoning abilities, greater contextual understanding, and improved consistency across generations.
$\bullet$ The initialization phase can be computationally intensive and less suitable for LLMs. As dataset size increases, LLM performance in this phase often degrades significantly, suggesting a need for auxiliary strategies or preprocessing.
$\bullet$ LLMs are sensitive to the volume and complexity of input data. When exposed to large-scale inputs, they are prone to generating infeasible or suboptimal solutions, highlighting the importance of integrated correction and repair mechanisms throughout the optimization process.
$\bullet$ Population-level LLM-based EVO offers greater computational efficiency compared to individual-level approaches. By operating on the entire population in a single prompt, it reduces the number of model calls and repetitive descriptions of operations.
$\bullet$ LLM-based reproduction is highly scalable in addressing network-structured problems with respect to token cost. This efficiency stems from the use of compact element ID representations, which scale logarithmically with dataset size, allowing LLMs to handle large problems more gracefully than if inputs scaled linearly.
II Related work
In this section, we will review the existing literature on the Synergy of LLM and optimization, and the application of evolutionary optimization on network-structured problems.
II-A Synergy of LLMs and optimization
Yu and Liu [27] comprehensively investigated the synergy of LLMs and evolutionary optimization and discussed different-angle applications. Wu et al. classified existing works regarding the synergy of LLMs and evolutionary computation into two main branches [10]: LLM-based black-box search operators [28] and LLM-based optimization algorithm automation [29, 30].
LLMs can potentially automate and optimize the design of evolutionary operators as indicated in [12], thereby reducing the need for manual tuning and domain-specific adjustments such as [31]. Yang et al. proposed to use LLMs as optimizer named Optimization by PROmpting (OPRO) to solve the traveling salesman problem [13], in which the previously generated solution and its evaluation value are used as part of the prompt for next generation. In the study by Meyerson [17], LLMs are employed as crossover operators to derive new solutions from parental inputs. Brownlee [32] also presented LLMs effectively functioning as mutation operators that enhance the search process. Liu et al. introduced a novel framework known as LLM-driven EA (LMEA) [14], which utilizes LLMs for both crossover and mutation operations. This approach highlights the adaptability of LLMs, where search behaviors can be easily modified by adjusting the LLM temperatures. Furthermore, LLM-based search operators can also be adapted for multi-objective scenarios by segmenting traditional optimization tasks into sub-problems [16]. In [33], Wang et al. explored the applicability of LLMs on constrained multiobjective optimization problems and achieved the promising results compared to the existing methods.
LLMs have shown significant potential in automated evolutionary optimization, autonomously optimizing, creating, and improving algorithms to effectively tackle optimization challenges [34]. As demonstrated in [19], heuristic searched by LLMs achieves excellent performance in different complex combinatorial problems. In [35], Liu et al. proposed a new method called Algorithm Evolution using the Large Language Model (AEL) that directly considers algorithms as individuals for evolution. AEL was then further extended to the design of guided local search algorithms [36], showing the strength of the LLM-based method over human-designed algorithms. After that, they extended AEL to an advanced model called Evolution of Heuristics (EoH) by exploring various prompts to solve different combinatorial tasks. LLMs analyze swarm intelligence algorithms to obtain a hybrid algorithm that combines various strengths of existing methods [37]. Mao et al. explored LLM-enhanced algorithm automation for identifying critical nodes [38]. In this approach, various heuristics are initialized as populations and then evolve with the assistance of LLMs. Morris et al. [29] introduced a thought reflection method to guide the LLMs-based algorithm automation and improved the accuracy of ExquisiteNetV2 models. The evolution process was also studied to enhance the adaptability and convergence by [39], where the mutation rate is adaptively adjusted with dynamic prompt.
LLMs can also be used as a surrogate model with the help of in-context learning to efficiently analyze the quality of the solution, as shown in [40]. Furthermore, LLMs can also assist in algorithm selection, as demonstrated in [41], through analyzing the code to grasp both its structural and semantic elements, along with the contextual understanding. An emerging direction worth noting involves using evolutionary optimization to search for the optimal prompt, enabling LLMs to achieve excellent performance. Notable examples of this approach can be found in the work of [42, 43, 44].
II-B Metaheuristic optimization on network-structured problems
Network-structured combinatorial problems play a crucial role in various practical domains due to their complexity and widespread applicability [45, 46]. The utility of evolutionary optimization in addressing discrete and non-linear problems has significantly facilitated its adoption across diverse fields, particularly in the context of complex networks [47, 48] and various combinatorial tasks [49, 50]. These optimization algorithms excel due to their inherent capability to navigate complex solution spaces effectively, making them suitable for tasks such as truck scheduling [51, 52] and job shop scheduling [53, 54]. In complex network, evolutionary optimization addresses several intricate combinatorial challenges, including influence maximization [55, 56], robustness analysis [57, 58], and community deception [59, 60]. Such tasks pose particular difficulties due to their non-linear interactions and discrete search spaces, factors that evolutionary algorithms are particularly adept at handling. For instance, Zhang et al. leveraged multi-objective evolutionary optimization to efficiently identify local communities within complex networks. Similarly, Liu et al. [61] presented an evolutionary framework employing a relationship-driven strategy to identify critical node subsets, achieving significant improvements across multiple benchmarks. Notably, evolutionary algorithms have also been effectively applied in network-related tasks such as important node identification [62], community discovery [63, 64], network reconstruction [65, 66], and network module recognition [67, 68]. These applications underscore the versatility and robustness of evolutionary optimization methodologies in addressing complex network problems.
A critical aspect of evolutionary optimization frameworks is the encoding scheme, which translates a problem into a suitable representation for algorithmic manipulation. While encoding schemes vary significantly based on the nature of the tasks, such as permutation-based encodings for scheduling tasks and binary or graph-based encodings for network tasks, the underlying mechanism of manipulating these representations remains consistent within the evolutionary algorithmic framework. Operations such as crossover and mutation are universally applied, regardless of the encoding specifics in the same way, i.e., exchange and replace elements. Thus, despite differences in representational strategies, evolutionary algorithms maintain a unified approach to exploring and exploiting the solution space.
III LLM-based evolutionary optimizer
@@
In the combinatorial problems in complex networks, the solution representation within the evolutionary optimization framework is generally defined as
@@
where X depends on the specific task and $n$ is the solution size [69]. In this work, we choose the influence maximization [70, 71] as an illustrative example. The problem is defined as: Given a graph $G=(V,E)$, where $V$ represents the set of nodes and $E$ represents the set of edges, the objective is to find a subset of nodes $S\subseteq V$ that maximizes the influence across the network. As for network-related problems, the solution representation can simply be
@@
Note that our focus is not on solving any specific network-related problem but on evaluating the performance of LLMs as operators for combinatorial problems in complex networks. The reason for selecting this problem is that it is general enough to provide insights and observations for LLMs as EVO due to the representation similarity on combinatorial problems. To enhance generality, ‘element’ will replace ‘node’, and ‘dataset’ will replace ‘graph’ in the rest of the discussion. For the selection, the input to LLMs is
@@
The first phase, Initialization is defined as $\mathcal{P}_{0}=\mathbf{F}_{I}(G)$, where $\mathcal{P}_{0}$ refers to the initialized population and $G$ refers to the input data. A strategic reduction in search space can speed up convergence and improve effectiveness [72]. However, based on current research [21], it is not feasible to feed datasets directly to LLMs, either in the form of adjacency matrices or text descriptions, because the reasoning ability of current LLMs on graphs remains limited. Therefore, instead of providing LLMs with just a dataset, we opt to input element IDs along with easily accessible metrics. Following this, LLMs will be instructed to (1) Rank the elements based on a specific metric; (2) Select a percentage of top elements as candidates; and (3) Sample filtered elements to form the initialized population. Several basic restrictions must be followed to ensure the quality of the initialized population: (1) Each solution in the population should contain the same number of elements with no duplicates; (2) The solutions within the population should be as diverse as possible; and (3) elements outside the candidate set are not allowed to appear in the population.
@@
The four components, initialization, selection, crossover, and mutation are all achieved by LLMs. The pseudocode of population- and individual-level LLM-based EVO is shown in Algorithms 1 and 2 respectively. The key difference is in line 7 of Algorithm 2, where solutions from the population are iteratively selected for reproduction, unlike in population-level optimization, which does not require this step. This generality is also reflected in our prompt engineering where little network domain knowledge is needed (please refer to Tables III and IV for LLM-based mutation).
Algorithm 1 Population-level LLM-based EVO with repair mechanism
Input: Dataset $G$, fitness function $f$, maximum number of generations $N_{\text{max}}$
Output: Optimized node set for influence maximization
1:Initialize population $\mathcal{P}=\{\mathcal{S}_{1},\mathcal{S}_{2},\cdots,\mathcal{S}_{\mathcal{N}_{P}}\}$ instructed by LLMs
2:Check and (repair) the output
3:while iteration count $<N_{\text{max}}$ do
4:     Calculate fitness value of solutions in $\mathcal{P}$: $f(\mathcal{P})$
5:     Select solutions as $\mathcal{P}$ for reproduction instructed by LLMs
6:     Check and (repair) the output
7:     Perform crossover on population instructed by LLMs
8:     Check and (repair) the output
9:     Perform mutation on $\mathcal{P}$ instructed by LLMs
10:     Check and (repair) the output
11:end while
12:Return solution with the highest fitness value in the population.
Algorithm 2 Individual-level LLM-based EVO with repair mechanism
Input: Dataset $G$, fitness function $f$, maximum number of generations $N_{\text{max}}$
Output: Optimized node set for influence maximization
1:Initialize population $\mathcal{P}=\{\mathcal{S}_{1},\mathcal{S}_{2},\cdots,\mathcal{S}_{n}\}$ instructed by LLMs
2:Check and (repair) the output
3:while iteration count $<N_{\text{max}}$ do
4:     Calculate fitness value of solutions in $\mathcal{P}$: $f(\mathcal{P})$
5:     Select solutions as the population for reproduction instructed by LLMs
6:     Check and (repair) the output
7:     for $S\in\mathcal{P}$ do
8:         Perform crossover on $S$ instructed by LLMs
9:         Check and (repair) the output
10:         Perform mutation on $S$ instructed by LLMs
11:         Check and (repair) the output
12:     end for
13:end while
14:Return solution with highest fitness value in the population.
III-A Computational cost of population- and individual-level LLM-based EVO
@@
Crossover: Let $\mathcal{V}$ be the token cost of LLMs, and $\mathcal{V}(P)$ and $\mathcal{V}(S)$ refer to the token of the entire population and individual solution, respectively. The input to LLMs consists of prompt $\mathcal{T}_{C}^{S}$ ($\mathcal{T}_{C}^{P}$) and a pair of solutions (population), where $\mathcal{V}(\mathcal{T}_{C}^{S})\approx\mathcal{V}(\mathcal{T}_{C}^{P})$. Note that $\mathcal{V}(\mathcal{T}_{C}^{P})$ is slightly higher than $\mathcal{V}(\mathcal{T}_{C}^{S})$ because there will be an additional text like ”Please randomly select pairs of solutions and continue applying the crossover operation until the number of newly created solutions matches the population size.” in the population-level method, in which the token cost is $\mathcal{V}(\mathcal{T}_{C}^{P})+\mathcal{V}(P)$. In contrast, the individual-level method requires LLMs to implement crossover operation on pair by pair, which costs $(\mathcal{V}(\mathcal{T}_{C}^{S})+2*\mathcal{V}(S))*\frac{\mathcal{N}_{p}}{2}$ where $\mathcal{N}_{p}$ refers to the population size. As $\mathcal{V}(P)=\mathcal{V}(S)*\mathcal{N}_{p}$, the cost difference between individual- and population-level crossover is $\mathcal{V}(\mathcal{T}_{C}^{S}*\mathcal{N}_{p}/2)-\mathcal{V}(\mathcal{T}_{C}^{P})\approx\frac{\mathcal{N}_{p}-2}{2}*\mathcal{V}(\mathcal{T}_{C}^{S})$.
@@
TABLE I: The errors of LLM outputs during evolutionary optimization.
III-B Error repair
@@
TABLE II: The required validations for the LLM output in different phases of evolutionary optimization.
The probabilistic nature of LLMs can lead to occasional undesirable results. As evolutionary optimization is an iterative process, an error early in the process can ‘blow up’ in subsequent cycles, creating a cascading effect to ruin the optimization process. To investigate the reliability of LLM-generated outputs across different phases, we establish rigorous validation standards and systematically investigate potential errors encountered during simulation. These errors are categorized into format errors and quality errors, with the latter further divided into critical and moderate types. Specifically,
@@
The summary of the different types of errors are listed in Table I and details are introduced as follows.
III-B1 Format Error
The format error refers to those that can directly interrupt the running of a program. Since format errors are rare and difficult to fix, it is usually more efficient to request a new generation.
  • E1: The output is not in the required format.
    • Example: When we ask LLMs to generate a list in Python, we expect only that list without any explanation and text for further operations. However, LLMs sometimes add extra text, like ”You are doing a crossover, and the result is …”.
    • Impact: This error can disrupt the optimization process because algorithms depend on a specific data structure.
  • E2: The output contains non-integer elements.
    • Example: The combinatorial problems require discrete integer values while LLMs sometimes output a solution containing non-integer numbers such as [3, 5.9, 2.1, 8].
    • Impact: Non-integer outputs can lead to invalid candidate solutions that cannot be evaluated, thus this error will also interrupt the process as E1.
III-B2 Critical Error
Most critical errors undermine diversity, which is essential for successful optimization. Note that once diversity is severely damaged, then it is very difficult to repair as only mutation contributes to exploration in the entire evolutionary optimization. Therefore, this kind of error is not allowed in the optimization. In a manner consistent with addressing format errors, a new output will be requested instead of attempting to correct an unacceptable one.
  • E3: The selected candidates significantly deviate from the ground truth.
    • Example: We input all possible elements, each with a specific metric, and seek the top 50% (for example). However, LLMs outputs may only minimally overlap with the ground truth, indicating that most of their suggestions are not suitable candidates.
    • Impact: This case will reduce the quality of the initialized population, further impacting subsequent optimization.
  • E4: The size of candidates falls significantly short of meeting the requirements.
    • Example: We input all possible elements, each with a specific metric, and seek the top 50%. However, LLMs may output only 10% elements.
    • Impact: Such a number is insufficient to guarantee the diversity of the population as the solutions in the initialized population will be very similar.
  • E5: The size of the population falls significantly short of meeting the requirements.
    • Example: We input the entire population consisting of 30 solutions into LLMs but only receive 10 solutions as output.
    • Impact: The drastic reduction in population size will severely affect the diversity of the population. The population size may be restored to the predefined number but the diversity remains poor as there will be a lot of repetitive solutions.
  • E6: The selected population contains one solution too many times.
    • Example: We input a set of solution IDs and wish to filter those low-fitness solutions. However, LLMs may occasionally produce repetitive IDs for numerous identical high-fitness solutions.
    • Impact: Although the low-fitness solutions are filtered, over-selecting a single solution damages the diversity of the population.
  • E7: Any solution appears in the population where all elements are the same.
    • Example: LLMs sometimes return several solutions like $[A_{1},A_{1},A_{1},A_{1}]$ but the constraint is that no repetitive element is allowed within the same solution.
    • Impact: When this kind of error occurs, we discovered that the number of erroneous solutions in the population is not a few but a lot. After several rounds of crossover, all solutions of the entire population will be polluted, resulting in poor diversity.
  • E8: The number of different elements in the population changes significantly.
    • Example: When we input a population containing 50 different elements, LLMs sometimes will output a population containing only 20 elements.
    • Impact: The drastic reduction in the number of different elements will severely affect the diversity of the population.
  • E9: The number of different solutions in the population changes significantly.
    • Example: This error is similar to E6 that occurs during selection, whereas this error occurs during reproduction. When we input a population consisting of 30 various solutions to LLMs, the output LLMs may also contain 30 solutions but most of which may be identical.
    • Impact: The drastic reduction in the number of different solutions will severely affect the diversity of the population.
III-B3 Moderate Error
The moderate errors have a high chance of being repaired during the optimization thus they will not yield severe impact.
  • E10: The size of some solutions fails to meet the requirement.
    • Example: When we input a solution consisting of 10 elements to LLMs for reproduction, the output may have 9 or 11 elements.
  • E11: The solution (population) after the operation remains the same as before.
    • Example: When we input a solution (population) to LLMs for reproduction, the output remains unchanged. For example, the solution $[A_{1},A_{2},A_{3},A_{4}]$ is still $[A_{1},A_{2},A_{3},A_{4}]$ after mutation.
  • E12: The size of the population fails to meet the requirement.
    • Example: When we input a population consisting of 30 solutions for reproduction, LLMs may output a population containing 28, 29, 31, or 32 solutions.
  • E13: The new solution contains duplicated elements.
    • Example: The issue resembles E7, but is less severe and involves only 2 or 3 identical elements in the output solution.
  • E14: The new solution contains invalid elements not found in candidate nodes.
    • Example: Suppose we have 100 elements and filter the 50 as candidates, the LLMs sometimes output a solution containing the element in the other 50. We found that LLMs did not produce elements outside these 100, thus we categorize this error as moderate.
  • E15: The selected population contains those with very low fitness.
    • Example: This error is like E11 but this is for selection. When we input a list of solutions to LLMs for selection, their output will be the same as what we input before.